On five tests, on which scores could range anywhere from 0 to 100, Juan had an average of exactly 88. What is the lowest score Juan
could have received on one test?
a) 0 b) 20 c) 22 d) 40 e) 44

Answer by nerdybill(7384)

On five tests, on which scores could range anywhere from 0 to 100, Juan had an average of exactly 88. What is the lowest score Juan
could have received on one test?
Assuming Juan has 100 on 4 out of 5 tests
then
Let x = Juan's lowest score
average:
(100 + 100 + 100 + 100 + x)/5 = 88
(400 + x)/5 = 88
400 + x = 440
x = 40
.
Answer: D